the hypotenuse of a right triangle is 13m long. if each leg were increased by 3m the hypotenuse would be increased by 4m. translate into equation.
Answer by nerdybill(7384)

the hypotenuse of a right triangle is 13m long. if each leg were increased by 3m the hypotenuse would be increased by 4m. translate into equation.
.
Let x = one leg
and y = second leg
.
since we have two unknowns, we'll need two equations.
.
x^2 + y^2 = 13^2
(x+3)^2 + (y+3)^2 = (13+4)^2